OpenStudy (anonymous):

Prove that the laplace transform of t^2 is 2/s^3

OpenStudy (anonymous):

well u need to evaluate this integral:\[\int_{0}^{\infty} t^2 e^{-st} dt\]it will be a integration by parts i guess :)

OpenStudy (kropot72):

\[L(t)=\int\limits_{0}^{\infty}e ^{-st}tdt\] \[=(t\frac{e ^{-st}}{-s})_{0}^{\infty}-\int\limits_{0}^{\infty}\frac{e ^{-st}}{-s}dt\] \[=\frac{1}{s ^{2}}\] since for s > 0 \[\lim_{t \rightarrow \infty}(te ^{-st})=0\] Now we have \[\frac{1}{s ^{2}}=\int\limits_{0}^{\infty}e ^{-st}tdt\] Differentiating both sides with respect to s \[-\frac{2}{s ^{3}}=\frac{d}{ds}\int\limits_{0}^{\infty}e ^{-st}tdt\] \[=\int\limits_{0}^{\infty}\frac{\partial}{\partial s}(e ^{-st})tdt\] \[=\int\limits_{0}^{\infty}-te ^{-st}tdt\] so that \[\frac{2}{s ^{3}}=\int\limits_{0}^{\infty}e ^{-st}t ^{2}dt\] which gives the Laplace transform of t^2